# Pierre Auger Collaboration

### v1

@@ III.4.2 Calibration using hybrid events
Figure 9: Energy-dependent resolution of the DNN (dashed red line) obtained by subtracting the FD resolution (dashed grey line) from the standard deviation of the FD and the SD $X_{\mathrm{max}}$ reconstruction (red line) in the hybrid data set.
@@ III.5 Systematic uncertainties
The systematic uncertainties of the $\langle X_{\mathrm{max}}\rangle$ measurement using the SD are shown in Fig. 10(a). The $X_{\mathrm{max}}$-scale uncertainty of the FD, as inherited by the DNN during the calibration using hybrid measurements to remove the dependence on hadronic interaction models, is depicted as a dash-dotted line. It contains uncertainties regarding the reconstruction, the atmosphere, and the calibration of the FD. Whereas the latter is independent of energy, the energy dependence is caused by the former two contributions. At low energies, reconstruction uncertainties of the FD dominate. These are surpassed by atmospheric uncertainties with increasing energy since more distant showers can be detected with correspondingly larger corrections for the light transmission between the shower and the detector. For more details on the FD uncertainty, we refer to Ref. [20]. Uncertainties from the SD are denoted as dashed region and comprise the remaining uncertainties of the detector aging ($<0.5$ $\mathrm{g\,cm^{-2}}$), diurnal variations (1 $\mathrm{g\,cm^{-2}}$) — since the FD calibration is performed at night — and the uncertainty on the calibration using hybrid events. The calibration uncertainty has two parts, the uncertainty of the definition of the absolute $X_{\mathrm{max}}$ scale, which is estimated to be ±$\pm 2~{}\mathrm{g\,cm^{-2}}$, and the energy dependence of the calibration. To estimate the energy-dependent uncertainty of the calibration, we compare the assumed constant calibration to a calibration function linear in $\log_{10}(E_{\mathrm{SD}}/\mathrm{eV})$ and use the observed differences as the upper and lower uncertainty on our calibration (compare Fig. 18(a) in the appendix).
@@ IV Investigation of the UHECR mass composition using the Surface Detector
In Fig. 11, we present the energy evolution of $\langle X_{\mathrm{max}}\rangle$ and $\sigma(X_{\mathrm{max}})$ as reconstructed using the DNN based on SD data (black circles) and as obtained using the standard FD reconstruction (open grey squares). The statistical uncertainties are estimated using bootstrapping and are shown as vertical lines, whereas the systematic uncertainties, discussed in Section III.5, are depicted as brackets. We do not show the systematic uncertainty for the measurement of $\langle X_{\mathrm{max}}\rangle$ using the FD as it is part of the SD uncertainty due to the cross-calibration we conducted. Since for $\sigma(X_{\mathrm{max}})$ no calibration is performed, we show the systematics for both measurements. The measured data are compared to predictions [66] for protons (red) and iron (blue) of the three hadronic interaction models EPOS-LHC, Sibyll2.3d, and QGSJetII-04, denoted by different line styles. In the right plot, we further show the number of events in each bin of the SD data, which is the same in both plots. The evolution of $\langle X_{\mathrm{max}}\rangle$ in Fig. 11(a) as a function of energy shows an excellent agreement between the SD and the FD measurements with only very small deviations that can be explained purely by statistics. This extends the $X_{\mathrm{max}}$ measurements to 100 EeV and confirms the transition from a lighter to a heavier composition with increasing energy, also reported in previous SD-based studies using the risetime of signals in the WCDs [21].
@@ IV.1 Discussion of breaks in the elongation rate
Note that a one-to-one agreement of breaks and structures generally, in the measurements of $\langle X_{\mathrm{max}}\rangle$ and $\sigma(X_{\mathrm{max}})$, is not to be expected since a change in the mean logarithmic mass does not need to coincide with a similar change in the measurement of $\sigma(X_{\mathrm{max}})$, i.e., the composition mixing [16]. It would rather reveal a characteristic structure of the composition. Interestingly, breaks at similar positions in the energy evolution of $X_{\mathrm{max}}$ and $\sigma(X_{\mathrm{max}})$ can be obtained when fitting a simplified astrophysical model using the FD $X_{\mathrm{max}}$ data and the Auger spectrum as measured by the SD (see Fig. 3 and Fig. 6 in Ref. [10]). A dedicated analysis focusing on the astrophysical interpretation and investigating the non-trivial interplay between the spectrum, $\langle X_{\mathrm{max}}\rangle$, and $\sigma(X_{\mathrm{max}})$, is ongoing and will be discussed in a future publication.
@@ IV.1.1 Crosscheck and comparison with the FD
The obtained model exhibiting a characteristic structure beyond a constant change in the mean logarithmic mass has to be consistent with the FD measurements. The comparison of the elongation model with the FD and SD is presented in Fig. 13. The model describes the FD data adequately with $\chi^{2}/\mathrm{ndf}=1.1$, demonstrating the consistency of the model with FD data.
Additionally, we tested the obtained model by investigating the fit with a different binning in energy and with different selections of core distance, zenith angle, and azimuth angle. We further studied the influence of seasonal, diurnal (day-night), and aging effects. No significant effects could be found for any of the tests. Next, we investigated the existence of the features before applying calibrations and without applying quality cuts and found a statistical significance level larger than $5\sigma$ using the model featuring three breaks. Since our selection removes approximately half of the events, we estimate that the expected median significance for identifying the breaks with a data set of similar statistics, and same quality as the full data set without any cuts, is $3.4\sigma$. The finding of a $4.6\sigma$ significance using our data selection confirms the expectation that the significance should increase with improving data quality.
@@
Rejecting a constant elongation rate using the two-break model is very stable and above a significance of $3.4\sigma$ for all scenarios. Applying instead of the FD calibration a correction of the SD $X_{\mathrm{max}}$ reconstruction based on the expected composition bias of the Auger mix using simulations (compare Fig. 4), a constant elongation rate can be rejected by more than $5\sigma$ assuming EPOS-LHC, as the first break is strongly pronounced. Therefore, we find a robust indication at a $4.4\sigma$ level for structures beyond a constant elongation rate. However, more statistics and/or a reduction in energy-dependent uncertainties are needed to confidently reject the 2-break model, i.e., to investigate the existence and nature of the third break.
@@ V Summary
In this work, we have presented a study of the UHECR mass composition based on the first two moments of the distribution of depth of maximum, $X_{\mathrm{max}}$, of air shower profiles using surface detector data of the Pierre Auger Observatory recorded between 2004 and 2018. With the use of deep learning, a novel reconstruction technique was developed, enabling for the first time a precise reconstruction of $X_{\mathrm{max}}$ using the recorded time-dependent SD signals on an event-by-event level. Our approach included cross-calibration with the complementary FD to remove mismatches between simulations and measured data and investigate systematic uncertainties, highlighting the importance of an independent data set for calibrating and validating machine learning algorithms. After cross-calibrating the method using fluorescence observations, we have studied the energy evolution of $\langle X_{\mathrm{max}}\rangle$ and $\sigma(X_{\mathrm{max}})$ from 3 EeV up to 100 EeV. Due to the superior duty cycle of the SD in comparison to the FD, the statistics for composition studies using $X_{\mathrm{max}}$ are increased by a factor of ten for energies above 5 EeV, enabling for the first time a measurement of $\sigma(X_{\mathrm{max}})$, sensitive to the composition mixing, beyond 50 EeV. We have found excellent agreement of the $\langle X_{\mathrm{max}}\rangle$ measurement with previous studies using the FD and confirm the transition of $\langle\ln\mathrm{A}\rangle$ from a lighter to a heavier composition. Furthermore, our $\sigma(X_{\mathrm{max}})$ measurement, which is independent of the FD calibration, agrees very well with previous studies using fluorescence telescopes. The finding of a decrease in the fluctuations with energy is confirmed, indicating a transition to a heavier and purer composition. The observation of very small fluctuations appears to exclude a large fraction of light nuclei at the highest energies.
@@
With the increase in statistics, we have found evidence at a level of $4.4\sigma$ for a characteristic structure in the evolution of the mass composition beyond a constant elongation rate, considering both statistical and systematic uncertainties. The model describing our data best features three breaks in the energy evolution of the composition and is compatible with the FD measurements. The locations of the identified breaks are found at energies similar to the ankle, instep, and suppression features identified in the UHECR energy spectrum. An interesting structure, while not statistically significant, is visible in $\sigma(X_{\mathrm{max}})$, which could suggest breaks at similar energies. However, more statistics and reduced systematic uncertainties are needed to study the nature of the identified breaks and, in particular, investigate the existence of the third break.
@@ Acknowledgments
Figure 16: Energy-dependent contributions of the SD systematic uncertainties for the measurement of $\langle X_{\mathrm{max}}\rangle$ after calibrating DNN to FD observations.

### v2

@@ III.4.2 Calibration using hybrid events
Figure 11: Summary of the SD systematic uncertainties for the measurement of $\langle X_{\mathrm{max}}\rangle$ after calibrating DNN to FD observations. Only the hybrid calibration shows a distinct energy dependence.
@@ III.5 Systematic uncertainties
The systematic uncertainties of the $\langle X_{\mathrm{max}}\rangle$ measurement using the SD are shown in Fig. 12(a). The $X_{\mathrm{max}}$-scale uncertainty of the FD, as inherited by the DNN during the calibration using hybrid measurements to remove the dependence on hadronic interaction models, is depicted as a dash-dotted line. It contains uncertainties regarding the reconstruction, the atmosphere, and the calibration of the FD. Whereas the latter is independent of energy, the energy dependence is caused by the former two contributions. At low energies, reconstruction uncertainties of the FD dominate. These are surpassed by atmospheric uncertainties with increasing energy since more distant showers can be detected with correspondingly larger corrections for the light transmission between the shower and the detector. For more details on the FD uncertainty, we refer to Ref. [62]. The uncertainties from the SD are denoted as a hatched region and are summarized in Fig. 11. They comprise the remaining uncertainties of the detector aging ($<0.5$ $\mathrm{g\,cm^{-2}}$), diurnal variations (1 $\mathrm{g\,cm^{-2}}$) — since the FD calibration is performed at night — and the uncertainty on the calibration using hybrid events. The calibration uncertainty has two parts, the uncertainty of the definition of the absolute $X_{\mathrm{max}}$ scale, which is estimated to be ±$\pm 2~{}\mathrm{g\,cm^{-2}}$, and the energy dependence of the calibration. To estimate the energy-dependent uncertainty of the calibration, we compare the assumed constant calibration to a calibration function linear in $\log_{10}(E_{\mathrm{SD}}/\mathrm{eV})$ and use the observed differences as the upper and lower uncertainty on our calibration (compare Fig. 19(a) in the appendix).
@@ IV Investigation of the UHECR mass composition using the Surface Detector
In Fig. 13, we present the energy evolution of $\langle X_{\mathrm{max}}\rangle$ and $\sigma(X_{\mathrm{max}})$ as reconstructed using the DNN based on SD data (black circles) and as obtained using the standard FD reconstruction (open grey squares). The statistical uncertainties are estimated using bootstrapping and are shown as vertical lines, whereas the systematic uncertainties, discussed in Section III.5, are depicted as brackets [^8] For the detailed table, see Table 4. We do not show the systematic uncertainty for the measurement of $\langle X_{\mathrm{max}}\rangle$ using the FD as it is part of the SD uncertainty due to the cross-calibration we conducted. Since for $\sigma(X_{\mathrm{max}})$ no calibration is performed, we show the systematics for both measurements. The measured data are compared to predictions [67] for protons (red) and iron (blue) of the three hadronic interaction models EPOS-LHC, Sibyll2.3d, and QGSJetII-04, denoted by different line styles. In the right plot, we further show the number of events in each bin of the SD data, which is the same in both plots. The evolution of $\langle X_{\mathrm{max}}\rangle$ in Fig. 13(a) as a function of energy shows an excellent agreement between the SD and the FD measurements with only very small deviations that can be explained purely by statistics. This extends the $X_{\mathrm{max}}$ measurements to 100 EeV and confirms the transition from a lighter to a heavier composition with increasing energy, also reported in previous SD-based studies using the risetime of signals in the WCDs [21].
@@ IV.1 Discussion of breaks in the elongation rate
[…]
@@ IV.1.1 Crosscheck and comparison with the FD
The obtained model exhibiting a characteristic structure beyond a constant change in the mean logarithmic mass has to be consistent with the FD measurements. The comparison of the elongation model with the FD and SD is presented in Fig. 15. The model describes the FD data adequately with $\chi^{2}/\mathrm{ndf}=1.1$, demonstrating the consistency of the model with FD data. Additionally, we tested the obtained model by performing the fits using a different binning in energy and by dividing the data into different subsets binned in core distance, zenith angle, azimuth angle, and detector age, as well as season (winter/summer) and daytime (day/night). In each of the subsets, the determined breaks agreed with uncertainties with the identified model, and no significant deviations could be found. Next, we investigated the existence of the features before applying calibrations and without applying quality cuts and found a statistical significance level larger than $5\sigma$ using the model featuring three breaks. Since our selection removes approximately half of the events, we estimated that the expected median significance for identifying the breaks with a data set with the quality of the full data set is $3.4\sigma$. In this study, we performed an energy-dependent re-sampling of the full data to account for the energy-dependent efficiency of the fiducial cut, i.e., to ensure in each energy bin similar statistics as the fiducial data set. The finding of a $4.6\sigma$ significance using our data selection confirms the expectation that the significance should increase with improving data quality.
@@
Rejecting a constant elongation rate using the two-break model is very stable and above a significance of $3.4\sigma$ for all scenarios. Applying instead of the FD calibration a correction of the SD $X_{\mathrm{max}}$ reconstruction based on the expected composition bias of the Auger mix using simulations (compare Fig. 5), a constant elongation rate can be rejected by more than $5\sigma$ assuming EPOS-LHC, as the first break is strongly pronounced. Therefore, we find a robust indication at a $4.4\sigma$ level for structures beyond a constant elongation rate. However, more statistics and/or a reduction in energy-dependent uncertainties are needed to confidently reject the two-break model, i.e., to investigate the existence and nature of the third break.
@@ V Summary
In this work, we have presented a study of the UHECR mass composition based on the first two moments of the distribution of depth of maximum, $X_{\mathrm{max}}$, of air shower profiles using surface detector data of the Pierre Auger Observatory recorded between 2004 and 2018. With the use of deep learning, a novel reconstruction technique was developed, enabling for the first time a precise reconstruction of $X_{\mathrm{max}}$ using the recorded time-dependent SD signals on an event-by-event level. Our approach included cross-calibration with the complementary FD to remove mismatches between simulations and measured data and investigate systematic uncertainties, highlighting the importance of an independent data set for calibrating and validating machine learning algorithms. After cross-calibrating the method using fluorescence observations, we have studied the energy evolution of $\langle X_{\mathrm{max}}\rangle$ and $\sigma(X_{\mathrm{max}})$ from 3 EeV up to 100 EeV. Due to the superior duty cycle of the SD in comparison to the FD, the statistics for composition studies using $X_{\mathrm{max}}$ are increased by a factor of ten [^10] A detailed comparison can be found in Table 5. for energies above 5 EeV, enabling for the first time a measurement of $\sigma(X_{\mathrm{max}})$, sensitive to the composition mixing, beyond 50 EeV. We have found excellent agreement of the $\langle X_{\mathrm{max}}\rangle$ measurement with previous studies using the FD and confirm the transition of $\langle\ln\mathrm{A}\rangle$ from a lighter to a heavier composition. Furthermore, our $\sigma(X_{\mathrm{max}})$ measurement, which is independent of the FD calibration, agrees very well with previous studies using fluorescence telescopes. The finding of a decrease in the fluctuations with energy is confirmed, indicating a transition to a heavier and purer composition. The observation of very small fluctuations appears to exclude a large fraction of light nuclei at the highest energies, further excluding the flux suppression to be caused by a pure proton beam interacting with the cosmic microwave background. However, this observation is insufficient to disentangle whether the suppression arises from the maximum injection energy at the sources, propagation effects, or a combination of both.
@@
With the increase in statistics, we have found evidence at a level of $4.4\sigma$ for a characteristic structure in the evolution of the mass composition beyond a constant elongation rate, considering both statistical and systematic uncertainties. The model describing our data best features three breaks in the energy evolution of the composition and is compatible with the FD measurements. The locations of the identified breaks are found at energies similar to the ankle, instep, and suppression features identified in the UHECR energy spectrum. While not statistically significant, in $\sigma(X_{\mathrm{max}})$, two plateaus are visible, where the fluctuations seem to stay constant, interestingly, at energies at which the elongation rate is closer to that of a pure composition. However, more statistics and reduced systematic uncertainties are needed to study the nature of the identified breaks and, in particular, investigate the existence of the third break.